OpenStudy (anonymous):

Can anyone solve the inequality w/Steps ? -1≤ -4≤ 16

OpenStudy (anonymous):

it is only an inequality

OpenStudy (anonymous):

is there a variable?

OpenStudy (sasogeek):

the inequality is a true statement, unless there's a variable, there's nothing wrong with it at the moment, there's nothing to solve

OpenStudy (anonymous):

Sorry ! -4m !!

OpenStudy (anonymous):

-1/4>m>4 greater than or equal to

OpenStudy (sasogeek):

solve it in 2 parts, \[-1\le-4m\]and\[-4m \le16\]solve for m

OpenStudy (sasogeek):

can you do that though?

OpenStudy (anonymous):

Yes I can .

OpenStudy (anonymous):

m ≤ -4 ?

OpenStudy (sasogeek):

nahh, when you divide an inequality by a negative number, the sign changes... \[-4m \le16\]\[m \ge -4\] and\[-1\le-4m\]\[1/4\ge m\] therefore, \[1/4\ge m \ge -4\]

OpenStudy (anonymous):

Ohhh , Mmkay yeah I forgot about the other half of it . I need to remember that ! Thanks though .

OpenStudy (sasogeek):

you're welcome :)